In his own words.
"I am starting a new automated experiment - beginning now. There was a little bit of a false start as pressure began increasing, but was programmed to be low. Found that I had unplugged the pressure control solenoid. Now plugged in and the pressure is headed to its control point. Hopefully smooth sailing from here.
The DAQ data file will be copied to the Data_DAQ folder in the following Google drive master folder for this experiment:
The experiment is a high temperature experiment, but it is not with a Parkhomov-like fuel. The fuel is about 0.35g of Ni foam that has been electroplated with Pd. The plated foam was supplied by Dennis Cravens. Separated from this foam in the end with a zirconia felt plug is about 0.05g of LiAlD4 which was supplied by Mathieu Valat. So, the predominant hydrogen isotope will be D2, though there will be some H2 from the OH that is attached to the LiAlD4. There are some clean-up heating and vacuum cycles, and then the experiment will be run in comparatively (to Parkhomov's experiments) at high pressure - as much as the LiAlD4 can generate in decomposition up to 90 PSIA. The temperature will top out at 1100°C and will last about 36 hours. This experiment also has a few cycles between 190°C and 350°C which Dennis identified as critical loading temperatures for the Pd. From a setup standpoint, the reactor looks just the same as it did in the previous experiment. Only the fuel, hydrogen isotope species, and the heat/pressure protocol have changed.
Other notables... the vacuum system has its leak repaired (I used Rectorseal #5 instead of teflon tape on the joints) and a timer has been setup to run the vacuum pump once per hour for 5 minutes. This should keep the vacuum reservoir below 10 Torr. The bug in the DAQ Labview code that caused the .csv files to be missing a character has been fixed. The gamma spectrometer software has been modified to produce a true 24 hour time stamp and include the offset from UT in the stamp.
UPDATE#01 - Experiment Ended, next experiment detailed
Bob Higgins' 'Cravens' experiment ended, no excess heat. Announcement of next experiment...
In Bob's own words
"The "Experiment_Cravens_20170421" is now complete. No excess heat was detected in this system with a fuel of electroplated Pd on Ni foam. It was amazing how close to a COP=1.0 was maintained through the entire protocol - probably within 0.2% which was <300mW error at the high temperature. If there was excess heat, it would have been detected.
The plots from participant "Can" are here, and the one labeled, "cravens-pdf-1492953525.pdf", is the plot for the whole run:
If you look at the plot that is the next to the bottom one in the .pdf file, it shows power across the run and also a plot for the power required for the calibration run. The fueled run input power is the yellow curve and the corresponding curve for no fuel (green) is calculated vs. temperature from the fit to the calibration data. Across the whole experiment, you can see how well the settled fueled run input power matches the calibration (no-fuel) power. This match corresponds to a COP==1.0 and the numerical match is closer than can be discerned from the graph.
The automation worked well. Still missing in the automation is the copying of the data files to the Google drive. Maybe I can have that automated by next experiment.
I am hoping to start the next experiment in about a week. It will be an AH-50 Ni experiment with the Ni powder etched in ultrasonic HCl. While still wet from rinsing in de-oxygenated water, I plan to freeze the powder and then dry in a heated vacuum. It will be run with 10% LiAlH4."